# Bridging Social Media and Search Engines: Dredge Words and the Detection of Unreliable Domains

### v3

@@ Introduction
On February 24, 2022, the day Russia began its invasion of Ukraine, Jacob Creech, a fringe Qanon conspiracy theorist, posted a series of unsubstantiated claims on Twitter. These tweets insinuated that the US had created COVID-19, and implied Russia’s invasion of Ukraine was actually being carried out to shut down US-funded “biolabs” in Ukraine and to prevent another global pandemic (League 2022). Hours after the tweet, the conspiratorial website InfoWars published an article promoting the tweet, crediting its author for uncovering an “ulterior motive theory” (League 2022). The tweet began circulating to other conspiratorial sites, and Google queries for “US biolabs” and “Ukraine Biolabs” spiked in the following days  [^1] https://trends.google.com/trends/explore?date=today“%205-y&geo=US&q=Ukraine“%20biolabs&hl=en. Conspiratorial sites were likely the only relevant search results until Snopes debunked the claim later that day (Evon 2022).
@@
In some ways, this is a success story of the current paradigm of combatting misinformation; fact-checkers responded rapidly to a quickly-spreading falsehood. In other ways, this highlights the challenges present in reactive misinformation interventions. Despite quick debunking by fact-checkers, the conspiracy was nonetheless echoed by established news sources like Fox News, and later amplified on the floor of the US Senate(League 2022). Researchers later found that a coordinated campaign of social media accounts boosted the narrative on Twitter (Alieva, Ng, and Carley 2022). As of 2024, despite the underlying claims having been debunked by many fact-checkers (Chappell and Yousef 2022; Parachini 2022), a search engine audit study found that Google, Bing, and Yandex all returned URLs on their first page of results that promote the “Ukraine Biolabs” conspiracy (Kuznetsova et al. 2024). The Ukraine Bioloab example highlights an important and understudied dimension of misinformation spread: the interaction between social media and search engines.
@@
An alternative to reactive fact-checking approaches are proactive algorithmic content moderation approaches. These can include modifying recommendation and ranking systems to decrease the reach and virality of unreliable information sources. In search engines, this can mean downranking articles from unreliable domains, and on social media, this can mean ranking posts containing unreliable domains lower in user newsfeeds. For proactive approaches to succeed, platforms need systems that can rapidly detect and discover unreliable domains. Both classification and detection are important as unreliable websites can, and do, deploy a number of tactics to avoid blacklists (Carragher, Williams, and Carley 2024a). In this work, we present a novel approach to the tasks of unreliable domain detection and discovery that leverages signals from both large-scale social media data and webgraph data.
@@
Directing users to search engines can be an effective tactic to spread misinformation as, since 2016, global surveys have consistently found that individuals report higher trust in search engines than traditional media (Barometer 2024; McDuling 2015). Thus by pointing users to search engines, either explicitly or implicitly, bad actors can create a false sense of content reliability. Little is understood about these pathways, and research has only recently begun exploring the ways social media users end up in data voids. The concepts of problematic queries (Golebiewski and Boyd 2019) and keyword signaling (Tripodi 2019) are highly related, but have only been explored in case studies. The recently-proposed concept of search directives provides a clear and observable path that could direct users to unreliable content on search engines. A search directive is content that is explicitly intended to prompt an online search, e.g., user $a$ directs user $b$ to ”look up Chemtrails on Google” (Robertson et al. 2023). In the context of data voids, search directives as a concept have two limitations: First, while search directives can direct users to unreliable content, they can also direct users to innocuous or helpful information, e.g., directing users to look up song lyrics or mathematical theorems. Second, as the search trends for “Ukraine Biolabs” demonstrate, users can be driven to unreliable content through search engines by querying phrases that appear without explicit search directives. To help fill these gaps, we propose the concept of dredge words—terms or keyphrases for which unreliable domains rank highly on search engines.
@@
Finally, we provide the first exploration of dredge words on social media. We incorporate a small set of dredge words into an unreliable domain discovery process in an attempt to mimic how social media users may transition from social platforms to the discovery of misinformation sources via search engines. Surprisingly, we find that dredge words frequently surface social media URLs in top Google Search Engine Result Page (SERP) positions—i.e., the websites returned when a user enters a query on Google Search. This suggests the existence of a bidirectional path that often leads back to social media. We show our heterogeneous model greatly outperforms competing systems in the top-k discovery of unlabeled unreliable websites. We publicly release the code, webgraph data collected for this project, a dataset of 3,939 dredge words for 46 unreliable news domains, and the resulting SERPs for each of the dredge word queries [^2] Artefacts will be publicly released following anonymous review period.. Our findings demonstrate that both direct and indirect paths to misinformation provide important signals that can be leveraged by researchers and platform designers working to mitigate the spread of misinformation across digital ecosystems.
@@
Figure 3: A tweet about “Indigo Children”.
@@ Methods
A label-stratified 80/10/10 split on labeled websites is used to create training, validation, and test sets. Nodes that did not have reliability labels—the majority of nodes in all networks—were masked during training. For our one-mode, heterogeneous, and homogeneous experiments we use a simple 2-layer graph neural network using GraphSAGE convolutions proposed in (Hamilton, Ying, and Leskovec 2017). We ran experiments using graph attention networks (Veličković et al. 2017) and heterogeneous graph transformers (Hu et al. 2020), but found marginal accuracy gains at the cost of longer training times. As our primary classification interest is evaluating the impact of context—i.e., signals from different sources—on domain credibility classification, rather than finding the best architecture, we elect to use homogeneous and heterogeneous GraphSAGE layers for each respective model. We include a single baseline that uses a heterogeneous graph transformer as well. Each model consists of a SAGEConv layer with dropout, with a hidden dimension of 512, followed by a ReLU activation and a second SAGEConv layer with a log softmax activation function. For models trained with user text, we included an additional linear layer to align input feature sizes. For each model, we train for a maximum of 1,000 epochs with early stopping based on validation loss and a patience of 50. We use an Adam optimizer with a starting learning rate of 1e-3 and a cosine annealing learning rate scheduler [^7] Implementation is available at ¡removed for anonymity¿. The heterogeneous GNN, the most complex network we consider, contains 3.2M parameters, and no model exceeded 5 minutes of training time to reach our early stopping convergence condition. All models in this paper were run or trained on a single NVIDIA GeForce RTX 3080.

### v4

@@ Introduction
On February 24, 2022, the day Russia began its invasion of Ukraine, Jacob Creech, a fringe QAnon conspiracy theorist, posted a series of unsubstantiated claims on Twitter. These tweets insinuated that the U.S. had created COVID-19 and implied that Russia’s invasion was actually an effort to shut down U.S.-funded “biolabs” in Ukraine and prevent another global pandemic (League 2022). Hours after these tweets, the conspiratorial website InfoWars published an article promoting them, crediting Creech for uncovering an ”ulterior motive theory” (League 2022). The tweets soon circulated to other conspiratorial sites, and search interest in “U.S. biolabs” and “Ukraine biolabs” spiked in the following days  [^1] https://trends.google.com/trends/explore?date=today“%205-y&geo=US&q=Ukraine“%20biolabs&hl=en. Conspiratorial sites were likely the predominant search results until Snopes debunked the claim later that day (Evon 2022).
@@
In some ways, this represents a success of the current misinformation response paradigm—fact-checkers acted swiftly to address a rapidly spreading falsehood. However, it also highlights the limitations of reactive misinformation interventions. Despite the swift debunking by fact-checkers, the conspiracy theory was still echoed by established news sources like Fox News and later amplified on the floor of the U.S. Senate (League 2022). Researchers later found that a coordinated network of social media accounts helped boost the narrative on Twitter (Alieva, Ng, and Carley 2022). As of 2024, even though fact-checkers have repeatedly debunked the underlying claims (Chappell and Yousef 2022; Parachini 2022), a search engine audit study found that Google, Bing, and Yandex still return first-page results that promote the “Ukraine biolabs” conspiracy theory (Kuznetsova et al. 2024). This example highlights an important and understudied aspect of misinformation: the interaction between social media and search engines.
@@
An alternative to reactive fact-checking is proactive algorithmic content moderation. This can involve modifying recommendation and ranking systems to reduce the reach and virality of unreliable information sources. In search engines, this might mean downranking articles from unreliable domains, while on social media, it could mean ranking posts containing unreliable links lower in users’ newsfeeds. For proactive approaches to work, platforms need systems that can identify unreliable domains. Both classification and discovery are essential as unreliable websites can and do employ tactics to evade blacklists (Carragher, Williams, and Carley 2024a). In this work, we present a novel approach to unreliable domain detection and discovery that leverages signals from both large-scale social media data and webgraph data.
@@
Little is understood about how users transition from social media to search engines to access unreliable websites. However, it is known that directing users to search engines can be an effective tactic for spreading misinformation. Since 2016, global surveys have consistently found that individuals trust search engines more than traditional media (Barometer 2024; McDuling 2015). Thus by directing users to search engines, either explicitly or implicitly, bad actors can foster a false sense of content reliability. Research has only recently begun exploring these pathways, and some helpful concepts like problematic queries (Golebiewski and Boyd 2019), data voids (Golebiewski and Boyd 2019), and keyword signaling (Tripodi 2019), have been examined in case studies. The recently proposed concept of search directives offers a clearer and more observable path by which users are directed to unreliable content. A search directive is content explicitly intended to prompt an online search, such as user $a$ telling user $b$ to ”look up Chemtrails on Google” (Robertson et al. 2023).
However, search directives have two key limitations as content moderation tools. First, while they can lead users to unreliable content, they can also direct them to neutral or beneficial information, such as song lyrics or mathematical theorems (Robertson et al. 2025). Second, as seen in the “Ukraine biolabs” case, users can be driven to unreliable content through search engines even without being explicitly told to search something. To overcome these limitations, we propose the concept of dredge words—terms or keyphrases for which unreliable domains rank highly in search results.
@@
Finally, we provide the first exploration of dredge words on social media. We incorporate a small set of dredge words into an unreliable domain discovery process in an attempt to mimic how social media users may transition from social platforms to the discovery of misinformation sources via search engines. Surprisingly, we find that dredge words frequently surface social media URLs in top Google Search Engine Result Page (SERP) positions—i.e., the websites returned when a user enters a query on Google Search. This suggests the existence of a bidirectional path that often leads back to social media. We show our heterogeneous model greatly outperforms competing systems in the top-k discovery of unlabeled unreliable websites. We show our heterogeneous model greatly outperforms competing systems in the top-k discovery of unlabeled unreliable websites. We publicly release the code, webgraph data collected for this project, a dataset of 3,939 dredge words for 46 unreliable news domains, and the resulting SERPs for each of the dredge word queries [^2] https://github.com/CASOS-IDeaS-CMU/DredgeWords. Our findings demonstrate that both direct and indirect paths to misinformation provide important signals that can be leveraged by researchers and platform designers working to mitigate the spread of misinformation across digital ecosystems.
@@
Figure 3: A truncated tweet about “Indigo Children”.
@@ Methods
A label-stratified 80/10/10 split on labeled websites is used to create training, validation, and test sets. Nodes that did not have reliability labels—the majority of nodes in all networks—were masked during training. For our one-mode, heterogeneous, and homogeneous experiments we use a simple 2-layer graph neural network using GraphSAGE convolutions proposed in (Hamilton, Ying, and Leskovec 2017). We ran experiments using graph attention networks (Veličković et al. 2017) and heterogeneous graph transformers (Hu et al. 2020), but found marginal accuracy gains at the cost of longer training times. As our primary classification interest is evaluating the impact of context—i.e., signals from different sources—on domain credibility classification, rather than finding the best architecture, we elect to use homogeneous and heterogeneous GraphSAGE layers for each respective model. We include a single baseline that uses a heterogeneous graph transformer as well. Each model consists of a SAGEConv layer with dropout, with a hidden dimension of 512, followed by a ReLU activation and a second SAGEConv layer with a log softmax activation function. For models trained with user text, we included an additional linear layer to align input feature sizes. For each model, we train for a maximum of 1,000 epochs with early stopping based on validation loss and a patience of 50. We use an Adam optimizer with a starting learning rate of 1e-3 and a cosine annealing learning rate scheduler [^7] Implementation is available at https://github.com/CASOS-IDeaS-CMU/DredgeWords. The heterogeneous GNN, the most complex network we consider, contains 3.2M parameters, and no model exceeded 5 minutes of training time to reach our early stopping convergence condition. All models in this paper were run or trained on a single NVIDIA GeForce RTX 3080.